Exercise · ZW-BINOMIAL-DISTRIBUTION-010 · ★★★ Advanced

A shooter hits with probability 0.8 on each shot, independently. In 3 shots, what is the probability of at least 2 hits?

Choose one option.

Show answer
  • B. \(0.896\)

Explanation

Add the cases of 2 and 3 hits: \(3(0.8)^2(0.2)+(0.8)^3=0.384+0.512=0.896\). The value \(0.384\) is only “exactly 2” and \(0.992\) is “at least 1”.

Study Probability and statistics in ZWMath →

Review the concept: Binomial distribution

A binomial variable \(X\) counts the successes \(k\) in \(n\) independent Bernoulli trials with the same success probability \(p\); \(P(X=k)\) is the probability of getting \(k\) successes.

\[P(X=k)=\binom{n}{k}p^{k}(1-p)^{n-k},\quad k=0,1,\dots,n\]

More practice problems: binomial distribution

← ZW-BINOMIAL-DISTRIBUTION-009